# A statistically consistent measure of Semantic Variability using Language Models

### v2

@@ Abstract
To address the challenge of variability in the output generated by language models, we introduce a measure of semantic variability that remains statistically consistent under mild assumptions. This measure, termed semantic spectral entropy, is an easily implementable algorithm that requires only standard, pre-trained language models. Our approach imposes minimal restrictions on the choice of language models, and through rigorous simulation studies, we demonstrate that this method can produce an accurate and reliable metric despite the inherent randomness in language model outputs.
A statistically consistent measure of Semantic Variability using Language Models
Yi Liu Seattle, Washington, USA
@@ 1 Introduction
.. The birth of Large Language Models (LLM) has given rise to the possibility of a wide range of industry applications Touvron et al. (2023); Chowdhery et al. (2023). One of the key applications of generative models that has garnered significant interest is the development of specialized chatbots with domain-specific expertise such as legal and healthcare Lexis ; Mesko (2023). These applications illustrate how generative models can improve decision-making and improve the efficiency of professional services in specialized fields.
This new LLM capability is made possible by the strong understanding of generative capabilities of the models Liu et al. (2023); Long (2023) and the advent of Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Gao et al. (2023). In an RAG system, the user interacts by submitting queries, which trigger a search for relevant documents within a pre-established database. These pertinent documents are retrieved based on the query and serve as a context for the LLM to generate an appropriate response. Since the implementation of RAG does not require a custom-trained LLM, it offers a cost-effective solution. The resulting chatbot can perform tasks traditionally handled by domain experts, improving operational efficiency and driving cost reductions.
However, a critical challenge impeding the widespread deployment of generative models in industry is the inherent variability present in these models Amodei et al. (2016); Hendrycks et al. (2021). Although parameters such as temperature, top-k, top-p, and repetition penalty are known to significantly influence model performance Wang et al. (2020, 2023); Song et al. (2024), even when these parameters are tuned to achieve deterministic output (e.g. setting temperature to 0 or top-p to 1), differences in the generated results can still occur in multiple runs. This persistent variability poses a significant barrier to the reliable and consistent application of generative models in practical settings.
Atil et al. (2024) conducted a series of experiments involving six deterministically configured large language models (LLMs), with temperature set to 0 and top p set to 1, across eight common tasks and five identical trials per task. The study aimed to assess the repeatability of model outputs by examining whether the generated strings were consistent between runs. The authors found that none of the LLMs demonstrated consistent performance in terms of generating identical outputs on all tasks Atil et al. (2024). However, the authors noted that when accounting for syntactical variations, the observed differences were relatively minor as many of the generated strings were semantically equivalent.
The variability in output has been attributed to the use of GPUs in large language model (LLM) inference processes, where premature rounding during computations can lead to discrepancies Nvidia (2024); Atil et al. (2024). Given this, it is reasonable to conclude that complete elimination of variability is unfeasible in any empirical setting. Consequently, we must acknowledge that the output of LLMs is inherently uncertain. In light of this, it becomes essential, similar to practices in statistics, to assess and quantify the level of uncertainty in the text generated by LLMs for any given scenario.
@@
Our work follows from a line of work inline with the concept of semantic entropy proposed in Kuhn et al. (2023); Nikitin et al. (2024); Duan et al. (2024); Lin et al. (2023). Kuhn et al. (2023) explore the entropy of the generated text by assigning semantic equivalence to the pairs of text and subsequently estimating the entropy. Similarly, Nikitin et al. (2024) and Lin et al. (2023) utilize graphical spectral analysis to enhance empirical results. However, a notable limitation in the entropy estimators proposed by Kuhn et al. (2023) and Nikitin et al. (2024) is their reliance on token likelihoods when assessing semantic equivalence, which may not always be accessible. Furthermore, Kuhn et al. (2023) acknowledge that the clustering process employed in their framework is susceptible to the order of comparisons, introducing variability into the results.
Moreover, previous work focuses on the empirical performance of the estimator. As such, while these methods have demonstrated favorable empirical outcomes, to the best of our knowledge, no authors have established using a theoretical analysis that their entropy estimators converge to a true entropy value as the sample size increases under an underlying generative model. Exploring the theoretical properties allows us to have a clear understanding of how the number of clusters and size of data would affect the estimator.
Our approach seeks to address these limitations by developing a robust theoretical analysis of the clustering procedure, ensuring convergence properties, and mitigating the variability inherent in prior methodologies. We propose a theoretically analyzable metric for quantifying the variation within a collection of texts, which we refer to as semantic spectral entropy. This measure addresses the observation that many generated strings, while lexically and syntactically distinct, may convey equivalent semantic content. To identify these semantic equivalences, we advocate the use of off-the-shelf generative language models (LMs). Moreover, we acknowledge that the LM used to evaluate semantic similarity is itself a stochastic generator. In response, we employ the well-established technique of spectral clustering, which is provably consistent under minimal assumptions on the generator, thereby ensuring the robustness and reliability of the proposed metric. Specifically, we demonstrate that the measure is statistically consistent under a weak assumption on the LM. To the best of our knowledge, this is the first semantic variability measure with proven convergence properties. As an empirical evaluation studies, we also propose a simple method for constructing clusters of different lexically and syntactically distinct but semantically equivalent text using compound propositions from Wittgenstein (2023).
@@ 2.1 Semantic entropy
.. We begin with a collection of textual pieces $n$, denoted $\mathcal{T}=(t_{1},\cdots,t_{n})$. Unlike that in Kuhn et al. (2023), our assumption is that we have access only to $\mathcal{T}$. In fact, we do not require the existence of a generative model and is interested only in variability of the semantics in the text. To evaluate the semantic variability of these texts in the context of a specific use case, we propose a theoretically proven measure of semantic entropy which we named semantic spectral entropy.
A key reason for opting against the use of variance as a measure of variability is that computing variance requires the definition of a mean, which is challenging to establish for semantic distributions. Although it is possible to define an arbitrary reference point, such as a standard answer in a chatbot that answers questions, evaluating the variability with respect to such a reference introduces bias.
@@
This formulation captures the uncertainty or disorder associated with assigning a given text $t$ to one of the clusters. Consequently, it provides a quantitative measure of semantic variability that avoids the biases introduced by arbitrary reference points.
@@
The relation $t_{i}\sim t_{j}$ if "$t_{i}$ is true if and only if $t_{j}$ is true" is an equiva
lence relation.
In light of the fact that equivalence relations can be defined arbitrarily based on the needs of the user. We propose that the determination of equivalence relations, denoted as $\sim$, is performed through a LM that generates responses independently of the specific generation of terms $t_{1},\dots,t_{n}$. However, we do not assume that we have access to probability distribution of the tokens as proposed by Kuhn et al. (2023); Nikitin et al. (2024) which is not always available. Rather, we just require a generator LM which can generate a determination of this relationship. Therefore, this LM can be general generative language model with a crafted prompt which we will use in our simulation studies. The error in this LM will be removed in the spectral clustering algorithm at the later stage. By leveraging this LM, we can define a function $e:{\mathcal{T},\mathcal{T}}\rightarrow{0,1}$, which is formally expressed as follows:
@@
However, since the function relies on an LM, $e(t_{i},t_{j})$ can be viewed as a Bernoulli random variable, whose value is dependent on the terms $t_{i}$ and $t_{j}$. did not address this issue but instead offers adopting a very powerful entailment identification model which the authors trust to identify the equivalence relation perfectly. In contrast, we suggest modeling the outputs of the LM as a random graph with an underlying distribution. In this framework, $t_{i}$ and $t_{j}$ represent nodes, while $e(t_{i},t_{j})$ are random variables that indicate the presence of an edge between the two nodes. Specifically, when $t_{i}\sim t_{j}$, the edge existence is governed by the following probability distribution:
@@ 3.1 Performance of spectral clustering algorithms
The proof is presented in the Appendix section A.2.1. We begin by presenting the result of strong consistency for the spectral clustering algorithm.
@@
The proof is provided in the Appendix section A.2. This establishes strong consistency result that we aim to present. At the same time, we also want to show the finite sample properties of the estimator ^$\hat{\mathcal{E}}(\mathcal{T})$.
@@
The full proof is provided in the appendix section A.3. A brief outline of the proof is as follows: we begin by using the results from Lei and Rinaldo (2015), which establish the rate of convergence for the stochastic block model. Next, we relate the errors of the spectral clustering algorithm to the errors in the empirical entropy, using the lemma 3.1 to establish this connection.
@@
The proof of this result is provided in the Appendix section A.3.1.
@@ 3. C3:
In Lemma A.6 and A.7 of the appendix, we show that |¯$|\mathcal{E}-\bar{\mathcal{E}}|\geq h\left(\frac{1}{p_{\min}}\right)K\sqrt{\frac{1}{2n}\log\left(2Kn\right)}$ with probability at most $\frac{1}{n}$.
In Lemma A.8 of the Appendix, we show that setting $c_{2}=2K\left(1-\sqrt{\frac{2\log(nK)}{np_{\min}}}\right)p_{\min}$, we have $2Kn_{\min}/n<c_{2}$ with probability at most $\frac{1}{n}$.
@@ 4 Simulation and data studies
  • Drone Flying/ Pilot Aerial drones
  • jazzercise / aerobics
To generate a cluster of text from this set of hobbies, we begin by randomly selecting $M$ items from a total of $N$ items in the list to formulate the compound proportion. This selection process yields $\binom{N}{M}$ potential subset of hobbies and we know that two subsets of hobbies are the same as long as their elements are the same. Next, to create individual text samples $t_{i}$ within the group, we randomly permute the order of the $M$ selected elements in the subset. This permutation process generates $M!$ unique samples for each combination of hobbies. Finally, the hobbies are then placed in its permuted order in a sentence like that below.
@@
In order to prevent models to rely on sentence structure, a few of these sentences are being designed.
We utilize Microsoft Phi-3.5 Abdin et al. (2024), OpenAI GPT3.5-turbo Hurst et al. (2024), A21-Jamba 1.5 Mini Lieber et al. (2021), Cohere-command-r-08-2024 Ustun et al. (2024),Ministral-3B Jiang et al. (2023) and the Llama 3.2 70B model Dubey et al. (2024) as $e$. These models are lightweight, off-the-shelf language models that are cost-effective to deploy and exhibit efficiency in generating outputs, thereby off-setting the computational cost of determining sermantic relationships. The exact prompt used to generate the verdict is specified in Appendix C.
ratio 0.2,0.3,0.5 0.3,0.3,0.4 0.5,0.5
@@
LLAMA 0.36 0.49 0.44 0.34 0.43 0.46 0.30 0.27 0.26
x MINISTRAL 0.22 0.27 0.13 0.25 0.23 0.21 0.14 0.22 0.21
COHERE 0.04 0.02 0.06 0.02 0.03 0.00 0.00 0.00 0.00
A21 0.05 0.00 0.00 0.00 0.01 0.00 0.00 0.00 0.00
PHI 0.08 0.07 0.07 0.03 0.03 0.00 0.00 0.00 0.00
GPT 0.06 0.02 0.00 0.01 0.00 0.00 0.00 0.00 0.00
@@
Table 1: Average $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ over simulation 10 iterations. We have three different ratio value run over three different data sizes. For $e$, we use Microsoft Phi-3.5 Abdin et al. (2024), OpenAI GPT3.5-turbo Hurst et al. (2024), A21-Jamba 1.5 Mini Lieber et al. (2021), Cohere-command-r-08-2024 Ustun et al. (2024), Ministral-3B Jiang et al. (2023) and the Llama 3.2 70B model Dubey et al. (2024).
@@
Figure 1: A scatter plot of $p-q$ against $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$. The different colors represents different language models used as $e$: A21 in blue, Phi in Orange, GPT in Green, Cohere in Red, Llama is Purple and Ministral in Brown. We notice that there is clear phrase change point where for $p-q<0.4$, we have that $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ is very high most of the time, for $p-q>0.4$, $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ is small with occasional jumps that the theory predicts.
$e$ $p-q$ $p$ $q$
LLAMA 0.17 0.17 0.00
MINISTRAL 0.22 0.99 0.77
COHERE 0.55 0.61 0.05
A21 0.81 0.96 0.15
PHI 0.67 0.67 0.01
GPT 0.80 0.87 0.07
Table 2: $p$, $q$ and $p-q$. For $e$, we use Microsoft Phi-3.5 Abdin et al. (2024), OpenAI GPT3.5-turbo Hurst et al. (2024), A21-Jamba 1.5 Mini Lieber et al. (2021), Cohere-command-r-08-2024 Ustun et al. (2024), Ministral-3B Jiang et al. (2023) and the Llama 3.2 70B model Dubey et al. (2024).
We complete simulation studies for a ratio of (0.2,0.3,0.5), (0.3, 0.3,0.4), and (0.5,0.5) and a sample size of 30, 50, 70. The average $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ over 10 iterations using different models as $e$ is recorded in table 1. The performance of algorithm using Cohere, A21, Phi, and GPT is strong while the performance of the algorithm with Minstral and Llama is weak. We primary attribute this to the inability of Llama and Minstral to make correct statements. $p-q$ is small for Llama and Minstral and large for Cohere, A21, Phi, and GPT (shown in Table 2). In fact, when we plot $p-q$ against $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ in Figure 1, we notice that there is phrase change at value $p-q=0.4$. $p-q<0.4$ $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ is high but $p-q>0.4$ implies that $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ is generally small. This phrase change is not predicted in the theory and suggests that more work is needed.

### v3

# A statistically consistent measure of semantic uncertainty using Language Models

@@ Abstract
To address the challenge of quantifying uncertainty in the outputs generated by language models, we propose a novel measure of semantic uncertainty, semantic spectral entropy, that is statistically consistent under mild assumptions. This measure is implemented through a straightforward algorithm that relies solely on standard, pretrained language models, without requiring access to the internal generation process. Our approach imposes minimal constraints on the choice of language models, making it broadly applicable across different architectures and settings. Through comprehensive simulation studies, we demonstrate that the proposed method yields an accurate and robust estimate of semantic uncertainty, even in the presence of the inherent randomness characteristic of generative language model outputs.
[…]
Appendix
A statistically consistent measure of semantic uncertainty using Language Models
Yi Liu Prorata.ai Bellevue, Washington, USA yi@prorata.ai
@@ 1 Introduction
.. The birth of large language models (LLMs) has given rise to the possibility of a wide range of industry applications Touvron et al. (2023); Chowdhery et al. (2023). One of the key applications of generative models that has garnered significant interest is the development of specialized chatbots with domain-specific expertise such as legal and healthcare. These applications illustrate how generative models can improve decision-making and improve the efficiency of professional services in specialized fields.
A significant challenge hindering the widespread industrial deployment of generative models is the lack of consistency in LLM outputs across multiple runs given the same input Amodei et al. (2016); Hendrycks et al. (2021). Without delving into the intricate mathematical architecture of generative language models, one can observe that these models function as stochastic generators: they take an input string and produce an output string. Prior studies Wang et al. (2020, 2023); Song et al. (2024) have demonstrated that the output for a fixed input can vary across runs, with the degree of variation influenced by factors such as temperature, top-k sampling, top-p sampling, and repetition penalties.
In a recent study, Atil et al. (2024) Atil et al. (2024) conducted experiments using six deterministically configured LLMs—each with temperature set to 0 and top-p set to 1—across eight common tasks, with five identical trials per task. The goal was to assess output repeatability under conditions designed to minimize randomness. Surprisingly, none of the models exhibited complete consistency in their outputs across all tasks. For more complex tasks, such as those involving college-level mathematics, the models frequently produced lexically distinct outputs across trials, resulting in zero consistency with respect to exact string matching. However, the authors also observed that many of these outputs were semantically equivalent despite their lexical differences, suggesting that syntactic variation does not necessarily imply divergence in meaning.
Variations in observed responses have been attributed to the use of GPUs during large language model (LLM) inference, where premature rounding and non-deterministic computation can introduce discrepancies Nvidia (2024); Atil et al. (2024). Consequently, it is reasonable to regard LLM outputs as inherently stochastic, with the elimination of this randomness being practically challenging—if not infeasible. This poses a significant concern for the commercial deployment of LLMs: when errors occur (i.e., when an output deviates from the expected result given a specific input), it becomes difficult to disentangle implementation-induced randomness from genuine model deficiencies. This ambiguity complicates efforts to make targeted improvements to the system. In response to this challenge, we propose the need for an empirical measure of uncertainty that evaluates the semantic variability of LLM outputs, rather than relying on surface-level lexical comparisons. Such a measure can be interpreted as a variance estimator that captures the semantic dispersion within a set of generated outputs.
@@
Our work follows from a line of work inline with the concept of semantic entropy proposed in Kuhn et al. (2023); Nikitin et al. (2024); Duan et al. (2024); Lin et al. (2023). Kuhn et al. (2023) explore the entropy of the generated text by assigning semantic equivalence to the pairs of text and subsequently estimating the entropy. Similarly, Nikitin et al. (2024) and Lin et al. (2023) utilize graphical spectral analysis to enhance empirical results. However, a notable limitation in the entropy estimators proposed by Kuhn et al. (2023) and Nikitin et al. (2024) is their reliance on token likelihoods when assessing semantic equivalence, which may not always be accessible. Furthermore, Kuhn et al. (2023) acknowledge that the clustering process employed in their framework is susceptible to the order of comparisons, introducing variation into the results.
Moreover, prior work in this area has primarily focused on the empirical performance of entropy estimators. While such methods have shown promising results in practice, to the best of our knowledge, no existing studies have provided a formal theoretical analysis demonstrating that these estimators converge to the true entropy value as the sample size increases under an assumed underlying model. Investigating the theoretical properties of our proposed measure enables a deeper understanding of how factors such as the number of clusters and the sample size influence the estimator’s behavior and consistency.
Our approach aims to address these limitations by developing a robust theoretical framework for the clustering procedure, ensuring convergence properties and reducing the variability observed in prior methodologies. We introduce a theoretically analyzable metric, semantic spectral entropy, to quantify the semantic variation within a collection of texts. This uncertainty quantification measure is motivated by the observation that many generated strings, though lexically and syntactically distinct, may express equivalent semantic content. To identify such semantic equivalences, we leverage standard generative language models (LMs) as tools for semantic evaluation.
Importantly, we recognize that the language model used to assess semantic similarity is itself imperfect, introducing potential noise due to its inability to provide uniformly accurate judgments across all text pairs. To mitigate this issue, we employ spectral clustering, a well-established technique known to be statistically consistent under mild assumptions on the underlying data-generating process. This provides robustness and theoretical guarantees for our proposed measure. Specifically, we show that semantic spectral entropy is statistically consistent under a weak assumption on the language model. To the best of our knowledge, this constitutes the first semantic uncertainty measure with formally proven convergence properties.
@@ 2.1 Semantic entropy
..We begin with a collection of $n$ textual items, denoted $\mathcal{T}=(t_{1},\ldots,t_{n})$. In contrast to the approach in Kuhn et al. (2023), we assume access only to the texts in $\mathcal{T}$, without requiring knowledge of the underlying generative process. While the methodology is designed with the goal of assessing uncertainty in outputs generated by large language models (LLMs), it is important to note that it does not necessitate $\mathcal{T}$ to be the result of a generative model. That is, the proposed method operates independently of the generation mechanism and can be applied in broader contexts. To evaluate the semantic uncertainty of the texts within a specific application domain, we introduce a theoretically grounded measure, which we term semantic spectral entropy.
While the proposed measure can be interpreted as a variance estimator for the semantics of $\mathcal{T}$, a central reason for not adopting the conventional notion of variance as a measure of uncertainty lies in the difficulty of defining a meaningful "mean" for semantic probability distributions. Unlike numerical data, semantic representations lack a natural central point around which variation can be measured. Although one might define an arbitrary reference point—such as a canonical or "standard" answer in the context of a chatbot—computing variance relative to such a reference introduces inherent bias and undermines the objectivity of the measure. Our approach, therefore, avoids reliance on predefined anchors and instead captures semantic dispersion directly, without the need for a fixed semantic center.
@@
This formulation captures the uncertainty or disorder associated with assigning a given text $t$ to one of the clusters. Consequently, it provides a quantitative measure of semantic uncertainty that avoids the biases introduced by arbitrary reference points.
@@
The relation $t_{i}\sim t_{j}$ if "$t_{i}$ is true if and only if $t_{j}$ is true" is an equivalence relation.
In light of the fact that equivalence relations can be defined arbitrarily based on the needs of the user. We propose that the determination of equivalence relations, denoted as $\sim$, is performed through a LM that generates responses independently of the specific generation of terms $t_{1},\dots,t_{n}$. However, we do not assume that we have access to probability distribution of the tokens as proposed by Kuhn et al. (2023); Nikitin et al. (2024) which is not always available. Rather, we just require a LM which can identify this relationship. In fact, this LM can be off-the-shelf language model with a crafted prompt, something we will use in our simulation studies. The error from LM will be removed in the spectral clustering algorithm at the later stage. By leveraging this LM, we define a function $e:{\mathcal{T},\mathcal{T}}\rightarrow{0,1}$, which is formally expressed as follows:
@@
However, since the function relies on an LM, $e(t_{i},t_{j})$ can be viewed as a Bernoulli random variable (given that we have established that there are inherent uncertainty in the LMs), whose value is dependent on the terms $t_{i}$ and $t_{j}$.Kuhn et al. (2023) did not address this issue but instead offers adopting a very powerful entailment identification model which the authors trust to identify the equivalence relation perfectly. In contrast, we suggest modeling the outputs of the LM as a random graph with an underlying distribution. In this framework, $t_{i}$ and $t_{j}$ represent nodes, while $e(t_{i},t_{j})$ are random variables that indicate the presence of an edge between the two nodes. Specifically, when $t_{i}\sim t_{j}$, the edge existence is governed by the following probability distribution:
@@ 3.1 Performance of spectral clustering algorithms
The proof is presented in the Appendix section B.2.1. We begin by presenting the result of strong consistency for the spectral clustering algorithm.
@@
The proof is provided in the Appendix section B.2. This establishes strong consistency result that we aim to present. At the same time, we also want to show the finite sample properties of the estimator ^$\hat{\mathcal{E}}(\mathcal{T})$.
@@
The full proof is provided in the appendix section B.3. A brief outline of the proof is as follows: we begin by using the results from Lei and Rinaldo (2015), which establish the rate of convergence for the stochastic block model. Next, we relate the errors of the spectral clustering algorithm to the errors in the empirical entropy, using the lemma 3.1 to establish this connection.
@@
The proof of this result is provided in the Appendix section B.3.1.
@@ 3. C3:
In Lemma B.6 and B.7 of the appendix, we show that |¯$|\mathcal{E}-\bar{\mathcal{E}}|\geq h\left(\frac{1}{p_{\min}}\right)K\sqrt{\frac{1}{2n}\log\left(2Kn\right)}$ with probability at most $\frac{1}{n}$.
In Lemma B.8 of the Appendix, we show that setting $c_{2}=2K\left(1-\sqrt{\frac{2\log(nK)}{np_{\min}}}\right)p_{\min}$, we have $2Kn_{\min}/n<c_{2}$ with probability at most $\frac{1}{n}$.
@@ 4 Simulation and data studies
To generate a cluster of text from this set of hobbies, we begin by randomly selecting $M$ items from a total of $N$ items in the list to formulate the compound proportion. This selection process yields $\binom{N}{M}$ potential subset of hobbies and we know that two subsets of hobbies are the same as long as their elements are the same. Next, to create individual text samples $t_{i}$ within the group, we randomly permute the order of the $M$ selected elements in the subset. This permutation process generates $M!$ unique samples for each combination of hobbies. Finally, the hobbies are placed in their permuted order in a sentence like that in the following.
@@
In order to prevent models from relying on sentence structure, some of these sentences are being designed.
We replicate this simulation setup in two different settings. In the first setting (Experiment 1), we consider ten common hobbies that a hypothetical individual engages in during their spare time. In the second setting (Experiment 2), we examine ten historical events that occurred on December 3, which we collect from Wikipedia Wikipedia contributors (2024). These two settings allow us to analyze the impact of text length by comparing long and short text scenarios.
We utilize Microsoft Phi-3.5 Abdin et al. (2024), OpenAI GPT3.5-turbo (denoted as GPT) Hurst et al. (2024), A21-Jamba 1.5 Mini (denoted as A21) Lieber et al. (2021), Cohere-command-r-08-2024 (denoted as Cohere) Ustun et al. (2024), Ministral-3B (denoted as Minstral) Jiang et al. (2023) and the Llama 3.2 70B model (denoted as Llama) Dubey et al. (2024) as $e$. These models are lightweight, off-the-shelf language models that are cost-effective to deploy and exhibit efficiency in generating outputs, thereby offsetting the computational cost of determining sermantic relationships. The exact prompt used to generate the verdict is specified in Appendix D.  [^1] Code are in https://github.com/yiliu9090/sermantic-spectral-entropy
Experiment 1 (Hobbies)
ratio 0.2,0.3,0.5 0.3,0.3,0.4 0.5,0.5 $p-q$ $p$ q
@@
LLAMA 0.36 0.49 0.44 0.34 0.43 0.46 0.30 0.27 0.26 0.17 0.17 0.00
MINISTRAL 0.22 0.27 0.13 0.25 0.23 0.21 0.14 0.22 0.21 0.22 0.99 0.77
COHERE 0.04 0.02 0.06 0.02 0.03 0.00 0.00 0.00 0.00 0.55 0.61 0.05
A21 0.05 0.00 0.00 0.00 0.01 0.00 0.00 0.00 0.00 0.81 0.96 0.15
PHI 0.08 0.07 0.07 0.03 0.03 0.00 0.00 0.00 0.00 0.67 0.67 0.01
GPT 0.06 0.02 0.00 0.01 0.00 0.00 0.00 0.00 0.00 0.80 0.87 0.07
Experiment 2 (Historical Events)
LLAMA 0.04 0.02 0.05 0.03 0.03 0.07 0.04 0.00 0.00 0.75 0.75 0.00
MINISTRAL 0.20 0.20 0.19 0.17 0.24 0.19 0.04 0.07 0.19 0.25 1.00 0.75
COHERE 0.08 0.04 0.09 0.08 0.09 0.04 0.00 0.01 0.04 0.52 0.98 0.46
A21 0.09 0.14 0.12 0.08 0.00 0.01 0.00 0.00 0.00 0.49 0.50 0.01
PHI 0.10 0.04 0.19 0.16 0.03 0.06 0.07 0.06 0.00 0.28 0.28 0.00
GPT 0.02 0.01 0.06 0.02 0.04 0.00 0.00 0.00 0.00 0.76 0.96 0.20
@@
Table 1: Average $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ over simulation 10 iterations. We have three different ratio value run over three different data sizes. For $e$, we use Microsoft Phi-3.5 Abdin et al. (2024), OpenAI GPT3.5-turbo Hurst et al. (2024), A21-Jamba 1.5 Mini Lieber et al. (2021), Cohere-command-r-08-2024 Ustun et al. (2024), Ministral-3B Jiang et al. (2023) and the Llama 3.2 70B model Dubey et al. (2024). We also present the average $p$, $q$ and $p-q$ observed in the experiments.
@@
Figure 1: A scatter plot of $p-q$ against $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ for outcome in experiment 1. The different colors represents different language models used as $e$: A21 in blue, Phi in Orange, GPT in Green, Cohere in Red, Llama is Purple and Ministral in Brown. We notice that there is clear phrase change point where for $p-q<0.4$, we have that $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ is very high most of the time, for $p-q>0.4$, $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ is small with occasional jumps that the theory predicts.
We conduct simulation studies for ratio configurations of (0.2, 0.3, 0.5), (0.3, 0.3, 0.4), and (0.5, 0.5), with sample sizes of 30, 50, and 70. The average deviation, $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$, over 10 iterations using different models as $e$ is recorded in Table 1.
In Experiment 1, the algorithm performs strongly when using Cohere, A21, Phi, and GPT, whereas its performance is weaker with Minstral and Llama. In Experiment 2, Llama, Cohere, A21, and GPT demonstrate strong performance, while Minstral and Phi exhibit weaker results. These findings suggest that Llama performs more effectively with longer text, which we primarily attribute to differences in $p-q$. Specifically, in Experiment 1, $p-q$ is small for Llama and Minstral but large for Cohere, A21, Phi, and GPT. In Experiment 2, $p-q$ is small for Phi and Minstral but large for Llama, A21, Cohere, and GPT.
A comparative analysis of Llama and Minstral in Experiment 1, as well as Phi and Minstral in Experiment 2, indicates that neither $p$ nor $q$ alone has a strong effect on $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$. Instead, when plotting $p-q$ against $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ in Figure 1, we observe a phase change at $p-q=0.4$. Specifically, for $p-q<0.4$, $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ is relatively high, whereas for $p-q>0.4$, $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$ tends to be lower. Notably, the experimental setup—defined by the number of data points and cluster ratio—has a considerably weaker effect on $|\bar{\mathcal{E}}-\hat{\mathcal{E}}|$.